Stepper Update 9/9 Please Read!

Gary J Sosa gsosa at stanford.edu
Tue Sep 9 16:25:40 PDT 2008


  Hello ASML Users

   Here is the most recent status of the stepper as commented in Coral:

   The ASML FSE's did quite a bit of diagnostic testing and  
troubleshooting of the illumination system. The X/Y/Z unit for lamp  
positioning was replaced to fix the lamp drift problem.  The FSE's  
found contamination on the bottom lens due to resist outgassing. After  
cleaning, the uniformity is now 1.58% and intensity is 700+ mw/cm2.   
We do still see an energy shift in the tool. The E/0 is 68-70 mj as  
developed on SVG developers. We still need to do more work in this  
area to understand and correct the problem. The tool will be up with  
cauton. IT IS STRONGLY ADVISED THAT YOU DO A FOCUS EXPOSURE MATRIX TO  
DETERMINE BEST ENERGY FOR YOUR LAYER, AND UPDATE THE EXPOSURE IN YOUR  
JOBS AS NEEDED.

   Please make the extra effort to do this so that your time is not  
wasted as well as chemicals in the event that you have to rework your  
wafers because of poor results.

  Lastly, The tool will still be down during the day on  
Wednesday(9/10) and Thursday(9/11) while ASML does a quarterly PM.  
Please check coral for current tool status.

   Thanks... Your SNF Staff




More information about the asml mailing list